OpenStudy (rainbow_dashie):

Numbers p and q are whole numbers. Which statements are always true? Choose all answers that are correct.

OpenStudy (rainbow_dashie):

A. p + q is a whole number B. p – q is a whole number C. p • q is a whole number D. is a whole number

OpenStudy (rainbow_dashie):

d.p/q

OpenStudy (anonymous):

u dunno this one

OpenStudy (anonymous):

srry sis

OpenStudy (rainbow_dashie):

Its ok:(

OpenStudy (anonymous):

only a and c

OpenStudy (rainbow_dashie):

OMG thx

OpenStudy (rainbow_dashie):

Is the set of multiples of 6 closed under subtraction? Explain why or provide a counterexample if not.

OpenStudy (rainbow_dashie):

A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12.

OpenStudy (rainbow_dashie):

c. 6-6=0 d.6-18=-12

OpenStudy (anonymous):

example 5+7 is a whole number 5-7 is not a whole number 5*7 is a whole number 5/7 is not a whole number.

OpenStudy (rainbow_dashie):

so.... a d

OpenStudy (anonymous):

eyupies

OpenStudy (rainbow_dashie):

ya sure sis

OpenStudy (rainbow_dashie):

d

OpenStudy (rainbow_dashie):

brb

OpenStudy (anonymous):

A what about 18-6=? it is not equal to 6 B is true if the given set is of integers but not true if it is a set of natural numbers.